OpenStudy (anonymous):

Suppose ln(x^2 - 1) = 9 Express the equation in the exponential form e^A = B. Use the variable x as necessary in your answer

OpenStudy (anonymous):

ln(x^2-1)=9 -> x^2-1=e^9 and manipulate all you want from there

OpenStudy (anonymous):

yeah thats what i got so far but how do i solve for x?

OpenStudy (anonymous):

Isolate variable x : x² = 1 + e^9 = .. Can you solve now?

OpenStudy (anonymous):

x=sqrt(1+e^9) that is the exact answer

OpenStudy (anonymous):

@MCronin Hey, you've spilled the secret :O

OpenStudy (anonymous):

@Chlorophyll lol i didnt even notice it was a different poster i thought it was the asker lol i need sleep